Painting A painter wishes to mix paint worth $15 per gallon with paint worth $18 per gallon to make a 12-gallon mixture worth $16.75 per gallon. How many gallons of each should he mix?
Answer by ankor@dixie-net.com(22740)

A painter wishes to mix paint worth $15 per gallon with paint worth $18 per gallon to make a 12-gallon mixture worth $16.75 per gallon. How many gallons of each should he mix?
:
Let x = amt of $18 paint
Total is to be 12 gallons therefore:
(12-x) = amt of $15 paint
:
15(12-x) + 18x = 16.75(12)
:
180 - 15x + 18x = 201
:
-15x + 18x = 201 - 180
:
3x = 21
:
x = 21/3
:
x = 7 gal of $18 paint
Therefore
5 gal of $15 paint